OpenStudy (anonymous):

What is the slope of the line? y=2x-7 y=2x-7 Slope= [ ]

OpenStudy (jojo4eva):

the slope of this line is 2

OpenStudy (anonymous):

@iGreen can you check this?

OpenStudy (jojo4eva):

the slope intercept form of a line is y=mx+b. M= the slope and 2 is in that position thus making the slope of the line 2. trust me its right

OpenStudy (anonymous):

\[\text{The slope of the line : $y=ax+b$ is always $a$ .}\]

OpenStudy (anonymous):

ok thanks @jojo4eva and @Noura11

OpenStudy (igreen):

y = 2x - 7 It's in the form of y = mx + b, where m = slope.

OpenStudy (igreen):

@jojo4eva Don't give answers.

OpenStudy (jojo4eva):

no problem my apologies @iGreen i just thought it was a simple enough question